The half-life of a zero order reaction is 3.15 sec for an initial concentration (Al)of 0.500 M. What is the rate constant for this reaction in mol/Ls?
I want the answer very clear and correct please with mol/L.s. Thank you
Answer #1

Ao = 0.500 M

t1/2 = 3.15 sec

K = ?

formula for rate constant = k = Ao / 2 t1/2

k = 0.500 / 2 x 3.15

k = 0.0794 mol /L s ------------------> answer